OpenStudy (anonymous):

-6+6 sqrt(3)i Express in trigonometric form.

OpenStudy (dumbcow):

\[r(\cos \theta + i \sin \theta) = -6 +i6\sqrt{3}\] sin and cos must be less than 1 r must be 12 \[\rightarrow 12(\frac{1}{2} - i \frac{\sqrt{3}}{2})\] now solve \[\cos \theta = \frac{1}{2}\] \[\sin \theta = -\frac{\sqrt{3}}{2}\]

OpenStudy (dumbcow):

oh sorry i switched the neg sign cos = -1/2 sin = sqrt3/2

OpenStudy (anonymous):

kk thx

OpenStudy (kenljw):

-6+i 6 sqrt(3) = 6 (-1 + i sqrt(3)) 6 sqrt((-1)^2 + sqrt(3)^2) exp{i inverse tangent(-sqrt(3)] 12 exp[-i1.0472] or in degrees 6 sqrt(2) exp[-i60] This commonly called polar form A cos(x) + iB sin(x)=sqrt(A^2+B^2) exp{iinverse tangent(A/B)

OpenStudy (kenljw):

sorry inverse tangent B/A

OpenStudy (kenljw):

Sorry |A + iB| = sqrt(A^2 + B^2) a angle of inverse tangent B/A therefore sqrt(A^2+B^2) (cos (inverse tangent(B/A) + isin(inverse tangent(B/A) sqrt(A^2+B^2) EXP{i{inverse tangent(B/A)}]